How do you solve this equation?
75- x/8=x/7

Answer by stanbon(75887)

75- x/8=x/7
-----
75 = (x/7 + x/8)
-----
75 = (1/7 + 1/8)x
----
75 = [(8+7)/56]x
---------
x = (56*75)/15
---------
x = 56*5
x = 280
------------------------
Cheers,
Stan H.